OpenStudy (anonymous):

Which of the following is equivalent to log base25^5 ?

OpenStudy (anonymous):

\[\log _{25}5\]

OpenStudy (anonymous):

\[\log _{25}5= \large \log _{5^2}5=\frac 1 2 \log_5 5 = \frac 12 \]

OpenStudy (anonymous):

In this case option 4.

OpenStudy (anonymous):

\[\log_a b = \frac {\log b}{\log a} \]